OpenStudy (anonymous):

Algerbra 1b plz help i'll fan and give a medal

OpenStudy (anonymous):

For the Lesson 1: The answer is A. We must have one and only one unique value of x for every y. For example: {(1,2), (1,3)} this is not a function, since the domain 1 has different in every pair. But this {(1,2), (2,3)} this is a function. For Lesson 2: The answer is C. In this equation throw the '+7y' part to the other side. You will get this:\[5x \ge 22 - 7y\] puy '-28' in place of y, you will get \[5x \ge 50 => x \ge 10\] For Lesson 3: Lets call leptop numbers 'x' and desktops 'y'. From the article we understand that: \[x = y + 8\] And thir sum will give us 156. \[x + y = 156\] \[2y + 8 = 156\] \[y = 74\] \[x = 82\] Lesson 4: Lets simplify this inequality, shall we. Together 'y's and 'x's in one place: \[-3x > 5y - 2y + 9\] \[-x > y + 3\] Throw the 3 other side and and multipy all sides by (-1), you will get the answer A.

OpenStudy (anonymous):

Lesson 6: This is a simple equation where you put 83 in place of 'C': \[83 = 24.50t + 9.50 = 3\] Lesson 7: The ansfer is F. Since we are depended on the number of studets, and number of students doesn't depend on something, the answer is F.